# Hydrodynamic Limit with a Weierstrass-type result

### v1

@@ 1. Introduction
Understanding the macroscopic behaviour of interacting particle systems is a central theme in probability and statistical mechanics. This work studies the local density of particles in a discrete Markovian system, which, after a rescaling in space and time, converges to the solution of a particular differential equation. From a probabilistic viewpoint, this corresponds to a law of large numbers. This passage from the discrete to the continuum, or from the microscopic to the macroscopic, is known in the physics literature as the hydrodynamic limit. Within the landscape of interacting particle systems, gradient systems play an important role, as their local structure makes them mathematically very tractable. These are models in which the algebraic current can be expressed as the discrete gradient of a potential [15, II, Subsection 2.4], further simplifying hydrodynamical analysis when compared to non-gradient models. For example, for a variety of models both equilibrium and non-equilibrium fluctuations can be obtained, as opposed to non-gradient models where this remains largely an open problem. Moreover, for gradient systems one can explicitly compute the diffusivity, whereas in general it is given by the Green Kubo formula [15, II, Subsection 2.2]. A physical interpretation has also been proposed for these systems, pointing to a dynamical term in the description of the general diffusivity that vanishes if [15, II, Subsection 2.2] and only if [13] the model is of gradient type, thus representing a purely static behaviour. It is therefore surprising that relatively few gradient models are known, and that general results within this class are scarce. The present work is a first attempt to address the fundamental question of how flexible such models can be: given a target diffusion equation, can one construct a microscopic dynamics whose empirical density evolves accordingly? It is worth mentioning that generality of functions has been investigated in the context of boundary conditions [10], allowing for generic functions via prescribed boundary dynamics evolving at the scale of the current, but a bulk analysis in this direction, as in the present work, becomes in some aspects more challenging due to the requirement of maintaining the gradient property. Here we focus solely on the bulk dynamics, defining the process in the one-dimensional discrete torus.
@@
To provide some additional context regarding possible diffusivities, a classical example of linear diffusivity arises from the Symmetric Simple Exclusion Process [9], arguably the simplest dynamics for which much is known. In this model, particles perform nearest-neighbour random walks with no directional bias in a discrete lattice and subject to an exclusion rule that prevents multiple occupancy of sites, yielding a Markov process with conserved density. For non-linear diffusivities, results are more scarce. Polynomial diffusivities were derived in [5], in the form of the Porous Media Model (PMM), allowing for diffusion coefficients of the form $D(\rho)=\rho^{m}$, for positive integer values of $m$. In [6], the possibility of extending the dynamics to non-integer values of the parameter $m$ was investigated. This question was answered affirmatively for values $m\in(-1,1)$ via the so-called Interpolating Model, allowing for a microscopic, continuous description of the transition from slow to fast diffusion. This was a surprising result because, although the dynamics is completely discrete, it is possible to obtain irrational exponents. It is relevant to mention that in [4, 3] such diffusivities were derived from processes that are continuous in nature, associated not with particles but coupled oscillators, in general a much more evolved context and with difficult tools.
The work [6] opened the door to the derivation of non-polynomial diffusion coefficients. A first attempt at such a generalisation was developed in [2], where the authors extended the polynomial basis of models to a long-jumps setting and, via a Taylor-like expansion, derived non-polynomial diffusivities associated with fractional equations. That work also highlighted a shortcoming of the choice of basis as the PMM family, as a large class of functions proved unattainable due to the emergence of negative transition rates when starting from the PMM. More precisely, the PMM family acts as a polynomial basis, but is not robust enough to guarantee non-negativity of the rates associated with specific linear combinations in order to derive the key diffusivity $D(\rho)=\rho^{n}(1-\rho)^{k}$, for arbitrary positive integers $n$ and $k$.
@@
This modelling issue was addressed in [11], where the present author further generalised the dynamics while preserving the gradient property and combinatorial interpretability. This led to the so-called ”Bernstein model”, which acts as an analogue of the Bernstein polynomial basis. Linear combinations of this basis, in the same spirit as [6], were analysed in [12], allowing for the introduction of the complete PMM family, continuously parametrised and yielding the diffusivity $D(\rho)=\rho^{m}$ for any $m>1$, thus resolving the main open problem posed in [6]. The aforementioned works paved the way for the analysis carried out in the present text, culminating in a first Weierstrass-type result.
@@
The model introduced here is different in nature from those studied in [12], and requires the introduction of some novel arguments. The main strategy for proving the hydrodynamic limit, however, is analogous, as we follow the well established entropy method [7]. This choice of methodology is justified both by the possibility of relying on previous and convenient results, and by the fact that it prepares the ground for a future complete open-boundary analysis. With this in mind, in the next sections we present only the main model and novel results, directing the reader to appropriate references for specific steps of the proof. An exception is the proof of the “block estimates”, which we include for completeness but relegate to Appendix A. We highlight the need for a novel ”replacement lemma” – Lemma 3.1 –whose necessity comes from the non-uniform continuity of the microscopic potential associated with the gradient property. This constitutes the main difference with respect to [12].
@@
At a more technical level, and pointing to future work and open problems, the restriction to positive diffusivities is crucial, as it ensures that the model is irreducible. Our main model belongs to the class of non-cooperative models, for which mobile clusters exist by definition. Concretely, it is a nearest-neighbour, non-cooperative, symmetric exclusion process of gradient type with non-local constraints. Particularly, the hopping rates of our model depend on a “mesoscopic” part of the configuration. Mobile clusters play a central role in the study of non-irreducible systems, often associated with diffusivities that can attain trivial values. In the present setting, however, allowing diffusivities that attain zero would give rise to mobile clusters whose length depends on the size of the system, making them too large to be analysed via standard arguments. This is in contrast with the work [12], where their size is independent of the length of the system. The nature of these clusters brings the system closer to a cooperative model, even though cooperative models are known to be non-gradient [14]. In this sense, the present model (extended to functions that attain zero) may act as a bridge between cooperative and non-cooperative models, and we believe that it should be studied using non-gradient tools. This is left for future work.
@@
One may also question the extension to higher dimensions. This can be performed straightforwardly by letting the process evolve as a one-dimensional model along each coordinate direction, yielding a higher-dimensional hydrodynamic equation. A more challenging question concerns the derivation of fractional hydrodynamic equations associated with long-range models. This is currently unattainable, as it would require a long-range extension of the Bernstein model introduced in [11]. Unfortunately, the reasoning used to extend the PMM to long jumps in [1] cannot be applied while maintaining the gradient property. This raises the question of whether a different argument or a different basis, macroscopically equivalent to the Bernstein basis, may exist. Both possibilities lead to interesting combinatorial problems that we intend to investigate in future work.
@@ 2.1. The model
where the generator $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{L}^{\beta,\ell_{N}}$}} is defined, for each $f:\Omega_{N}\to\mathbb{R}$, through
@@
It is convenient to shorten the rates
@@
Because the function $\beta$ is positive, the process $\mathcal{L}^{\beta,\ell_{N}}$ is irreducible. Moreover, being also symmetric and of gradient type, as we shall see shortly, the Bernoulli product measures with constant profile are invariant measures. Because $\beta$ is normalized, $\mathbf{c}_{\beta,\ell_{N}}\leq 1$. We shall make no further assumptions on $\beta$ beyond its boundedness, positivity and continuity.
It is crucial to note that the process associated with $\mathcal{L}^{\beta,\ell_{N}}$ is a superposition of the elementary processes introduced in [11]. Precisely, for each $0\leq n\leq\ell_{N}$, consider the generator $\mathcal{B}^{n,\ell_{N}}$, given by replacing in (2.4) the constraint $\mathbf{c}_{\beta,\ell_{N}}$ by $\mathbf{b}_{n,\ell_{N}}$, with
@@
In [11, Proposition 2.7] it is proved that the generator $\mathcal{B}^{n,\ell_{N}}$ is associated with a gradient model, in the sense of [15, II, Subsection 2.4]. Thus, from the linear combination in (2.8), our main process inherits this property. This is the content of the next lemma, whose proof is direct given the expression in [11, Proposition 2.7].
@@
$\mathbf{c}_{\beta,\ell_{N}}\mathbf{e}_{0,1}-\mathbf{c}_{\beta,\ell_{N}}\mathbf{e}_{1,0}$ $=-\nabla\mathbf{H}_{\beta,\ell_{N}},$ (2.10)
@@
$\mathbf{g}_{\beta,\ell_{N}}(\eta)$ $=\frac{1}{\ell_{N}+1}\sum_{j=1}^{\ell_{N}}\sum_{i=1}^{j}\tau^{-i}\mathbf{J}_{\beta}^{j}(\eta),$ (2.12)
$\mathbf{J}_{\beta}^{j}(\eta)$ $=\beta(\langle\eta\rangle_{j})(\mathbf{e}_{0,1}(\eta)-\mathbf{e}_{1,0}(\eta)).$ (2.13)
It will be convenient to note an equivalent expression for $\mathbf{h}_{\beta,\ell_{N}}$. Let ${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{P}_{n}}}(A)$ be the collection of subsets of some set $A\subset\mathbb{T}_{N}$ with exactly $n$ elements. One can then express (see the last two displays in the proof of [11, Proposition 2.7] for the precise computations)
@@ 2.2. Topological setting
The uniqueness of solutions of the weak formulation in last definition is consequence of the regularity in (1). This is very straightforwardly shown via Oleinik’s method (see, for instance, [6, Lemma B.3]).
@@ 3.1. General scheme
The link between our process and the weak formulation (2.17) is given through Dynkin’s martingale (see [9, Appendix 1, Lemma 5.1]),
@@
One can show that the sequence of probability measures $(\mathbb{Q}_{N})_{N\in\mathbb{N}}$ is tight with respect to the Skorokhod topology of $\mathcal{D}\left([0,T],\mathcal{M}_{+}\right)$ by following [12, Subsection 3.2], which invokes Aldous’ conditions. As evidenced in [6, Proposition 3.1] and [12, Subsection 3.2], tightness for exclusion processes in the torus satisfying the gradient condition is consequence of the following: (i) $|\mathbf{r}_{\beta,\ell_{N}}|_{\infty}=o(N)$ ; (ii) $|\mathbf{h}_{\beta,\ell_{N}}|_{\infty}$ being uniformly bounded from above by a constant independent of $N$; (iii) and the estimate in Lemma 4.1 vanishing as $N\to\infty$. The former two are straightforwardly identified, therefore we refer the reader to the aforementioned references for the specific details to include them in the proof.
@@
Provided the aforementioned convergence by subsequences and absolutely continuity of the limit measure, one then argues that the density $\rho$ satisfies the notion of weak solution in Definition 2.5. This encompasses two results to be held $\mathbb{Q}$ almost-surely: (i) satisfaction of the weak formulation $\mathfrak{F}_{t}(\rho^{\text{ini}},\rho,G)=0$; and (ii) regularity, as in (1) in Definition 2.5. The latter is typically shown via the so-called ”energy estimate”. This is standard in the literature, and shown to hold for models with properties satisfied by the model in the present text. Concretely, proved the ”replacement lemmas” (here stated in Corollary A.2 and A.3), and the vanishing, as $N\to\infty$, of the remainder term in the gradient property, (2.12), as in Lemma 4.1, the regularity property (1) is consequence of $\mathbf{c}_{\beta,\ell_{N}}$ being uniformly bounded from above by a constant independent of $N$. For the precise computations, we refer the reader to [6, Section 5], where this was shown for the ”Interpolating process” in the slow-diffusion regime; or [12, Section 5 – Regime I]. The latter is closer, in terms of notation and exposition, to the present text. The next step is to characterize the limit $\mathbb{Q}$.
@@ 3.2. Characterization of the limiting points
where we recall that $\mathfrak{F}_{t}(\rho^{\rm ini},\rho,G)$ is given in (2.17). The next steps up to (3.14) are presented in detail in [12, Proof of Proposition 3.2], from the start of the proof up to equation (3.12) in the aforementioned work. Here, we only overview the approach and present the main novelties.
@@
  • After an application of Doob’s inequality, one can see that the first term vanishes in the limit $N\to+\infty$ by repeating the computations to show the first equality in [12, Equation (3.2)], where $|\mathbf{r}_{\beta,\ell_{N}}|_{\infty}=o(N)$ is required;
  • For the second, one applies [12, Proposition 4.3] and then Lemma 4.1;
@@
$\mathbf{h}_{\beta,\ell_{N}}(\eta)-\mathbf{h}_{\beta,L}(\eta)$ $=\sum_{m=0}^{M}\beta(\tfrac{m}{M})\bigg\{\frac{1}{\ell_{N}+1}\sum_{n=0}^{P_{\ell_{N}}(\eta)-1}\text{B}_{m,M}(\tfrac{n}{\ell_{N}})-\frac{1}{L+1}\sum_{n=0}^{P_{L}(\eta)-1}\text{B}_{m,M}(\tfrac{n}{L})\bigg\}$ (3.18)
@@
Applying [12, Proposition 4.3] and the triangle inequality, the limit above is a direct consequence of forthcoming Lemma 4.2.
@@
Next, we apply [12, Proposition 4.3] to each of the remaining probabilities and analyse them separately with the Replacement Lemmas in Section 4 (Corollaries A.2 and A.3), after an application of the triangle inequality in order to pass the summations over $m$ to outside of the expectation. Precisely, the terms associated with (3.28) and (3.27) are analysed with the upcoming Lemma A.2, with $L$ in the statement of the aforementioned lemma either equal to $1$ or an arbitrary integer $L>1$, respectively, and $w$ and $y$ chosen accordingly. To treat the term associated with (3.26), it is enough to apply the triangle inequality and then invoke Lemma A.3. In order to conclude the proof, we perform the sequence of limits $\limsup_{M\to+\infty}\limsup_{L\to+\infty}\limsup_{\epsilon\to 0}\limsup_{N\to+\infty}$.

### v2

@@ 1. Introduction
Understanding the macroscopic behaviour of interacting particle systems is a central theme in probability and statistical mechanics. This work studies the local density of particles in a discrete Markovian system, which, after a rescaling in space and time, converges to the solution of a particular differential equation. From a probabilistic viewpoint, this corresponds to a law of large numbers. This passage from the discrete to the continuum, or from the microscopic to the macroscopic, is known in the physics literature as the hydrodynamic limit. Within the landscape of interacting particle systems, gradient systems play an important role, as their local structure makes them mathematically very tractable. These are models in which the algebraic current can be expressed as the discrete gradient of a potential [21, II, Subsection 2.4], further simplifying hydrodynamical analysis when compared to non-gradient models. For example, for a variety of models both equilibrium and non-equilibrium fluctuations can be obtained [1, 10, 11, 9] as opposed to non-gradient models where this remains largely an open problem (see discussion in [6] and references therein). Moreover, in general the diffusion coefficient is described with Green Kubo formula [21, II, Subsection 2.2], and particularising for gradient systems one can compute it explicitly with the knowledge of the invariant measures. This is convenient in devising models associated with a target diffusion. A physical interpretation has also been proposed for these systems, pointing to a dynamical term in the description of the general diffusivity that vanishes if [21, II, Subsection 2.2] and only if [19] the model is of gradient type, thus representing a purely static behaviour. It is therefore surprising that few general results within this class are available. The present work is a first attempt to address the fundamental question of how flexible such models can be: given a target diffusion equation, can one construct a microscopic dynamics whose empirical density evolves accordingly? It is worth mentioning that generality of functions has been investigated in the context of boundary conditions [16], allowing for generic polynomial functions via prescribed boundary dynamics evolving at the scale of the current. An analysis of the dynamics in the bulk, as in the present work, becomes in some aspects more challenging due to the requirement of maintaining the gradient property. In this way, we shall restrict the studies to the one-dimensional discrete torus.
@@
To provide some additional context regarding possible diffusivities, a classical example of linear diffusivity arises from the Symmetric Simple Exclusion Process [15], arguably the simplest dynamics for which much is known. In this model, particles perform nearest-neighbour random walks with no directional bias in a discrete lattice and subject to an exclusion rule that prevents multiple occupancy of sites, yielding a Markov process with conserved density. Gradient models can encapsulate polynomial diffusivity, with the Porous Media Model (PMM) [8] acting as a toy model for diffusion coefficients of the form $D(\rho)=\rho^{m}$, for positive integer values of $m$. The dynamics is easily described: under the exclusion rule, a pair composed by a vacant and occupied sites exchange their occupation values with rate given by the number of boxes composed by $m$ aligned particles around the pair. The derivation of non-polynomial diffusion from exclusion type dynamics was addressed in [12], where it was performed an extension of the dynamics to non-integer values of the parameter $m$ via a generalised binomial expansion, successful for values $m\in(-1,1)$. This allows a microscopic description of the transition from slow to fast diffusion, continuously parametrised, by a collection of gradient models. It is relevant to mention that the two-parts works [7, 5] identify the diffusivity $D(\rho)=\rho^{m}$ for $m\geq 1$ for processes that are continuous in nature, associated with coupled oscillators and where other tools apply.
@@
The work [4] follows on the direction of investigating non-polynomial diffusion in more generality and with long-range interactions. The works [12, 4] evidence a deficiency in the porous media collection in generating particular diffusion, yielding possibly ”negative rates” by definition, that was then addressed in [18]. In it, the author introduces a collection of gradient dynamics modelling the Bernstein polynomial basis and generalising the PMM. Linear combinations in this basis, in the same spirit as [12], were analysed in [18], allowing for the introduction of the complete PMM family, continuously parametrised and yielding the diffusivity $D(\rho)=\rho^{m}$ for any $m>1$, thus resolving the main open problem posed in [12]. The aforementioned works paved the way for the analysis carried out in the present text, culminating in a first Weierstrass-type result.
@@
The model introduced here has critical differences from those studied in [18], and requires the introduction of some novel arguments. The main strategy for proving the hydrodynamic limit, however, is analogous, as we follow the well established entropy method [13]. This choice of methodology is justified both by the possibility of relying on previous and convenient results, and by the fact that it prepares the ground for a future complete open-boundary analysis. With this in mind, in the next sections we present only the main model and novel results, directing the reader to appropriate references for specific steps of the proof. We highlight the need for novel arguments due to the non-uniform continuity of the microscopic potential associated with the gradient property (see Lemmas 3.1 and 4.2 and their applications in the proof of Theorem 2.8). This is the main technical challenge with respect to [18].
At a more technical level, and pointing to future work and open problems, the restriction to positive diffusivities is crucial, as it ensures that the model is irreducible.
1.1. Open problems and future work
@@
Our main model is irreducible, and the hopping rates depend on a “mesoscopic” part of the configuration: letting $N>>1$ be the length of the system, the constraints are dependent of the information in a ball of radius $\ell_{N}+1=o(N)$, where $\ell_{N}\xrightarrow{}\infty$ as $N\xrightarrow{}\infty$. If the function $\beta$ were allowed to take the value zero, the resulting dynamics would belong to the class of non-cooperative [20] models, for which ”mobile clusters” exist. Mobile clusters are blocks of configurations that can move together and facilitate the transport of mass thorough the system. These objects play a central phenomenological and technical role in non-irreducible systems, associated with diffusivities that can attain trivial value for specific densities [8, 2, 12, 18, 3, 4]. In the present setting, however, allowing diffusivities that attain zero would give rise to mobile clusters whose length depends on the size of the system, making them too large to be analysed via standard arguments. This is in contrast with [18], where their size is independent of the length of the system and therefore of completely local nature. The nature of these clusters brings the system technically closer to a cooperative model, that are we known to be non-gradient [20]. In this sense, it may act as a bridge between cooperative and non-cooperative models. This is left for future work.
@@
We also identified a phase transition related with the range of interaction $2\ell_{N}+1$, when $\ell_{N}=O(N)$. The author is currently investigating the hydrodynamics of the resulting system in this regime, when in contact with reservoirs. One may also question the extension to higher dimensions. This can be performed straightforwardly by letting the process evolve as a one-dimensional model along each coordinate direction, yielding a higher-dimensional hydrodynamic equation. A challenging question concerns the derivation of fractional hydrodynamic equations, associated with long-range models. This would require a “long-jumps” extension of the nearest-neighbour dynamics in [17], for which the arguments in [3] violate the gradient property.
@@ 2.1. The model
The generator $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{L}^{\beta,\ell_{N}}$}} is defined, for each $f:\Omega_{N}\to\mathbb{R}$, through
@@
The dynamics is illustrated in Figure 1. It is convenient to shorten the rates
@@
By translation invariance, the constraint for an exchange at a node $\{x,x+1\}$, with generic $x\in\mathbb{T}_{N}$, is defined via translation, with rate coinciding with $\tau^{x}\mathbf{r}_{\beta,\ell_{N}}$
[…]
Figure 1. The rate for an exchange at the node $\{0,1\}$ (box in red) depends on the density of particles in the windows $W_{0},W_{1},\dots,W_{\ell_{N}}$: it is given by the average, over these windows, of $\beta$ evaluated at the empirical average in the corresponding window, $\langle\eta\rangle_{j}$. Precisely, the exchange realises with rate $\frac{1}{\ell_{N}+1}\sum_{j=0}^{\ell_{N}}\beta(\langle\eta\rangle_{j})\mathrm{d}t$.
@@
Because the function $\beta$ is positive, the process $\mathcal{L}^{\beta,\ell_{N}}$ is irreducible. Moreover, being also symmetric and of gradient type, as we shall see shortly, the Bernoulli product measures with constant profile are invariant measures. Because $\beta$ is normalized, $\mathbf{c}_{\beta,\ell_{N}}\leq 1$. We shall make no further assumptions on $\beta$ beyond:
  • boundedness;
  • positivity;
  • continuity.
It is crucial to note that the process associated with $\mathcal{L}^{\beta,\ell_{N}}$ is a superposition of the elementary processes introduced in [17] and representing the Bernstein polynomial basis. For each $0\leq n\leq\ell_{N}$, consider the generator $\mathcal{B}^{n,\ell_{N}}$, given by replacing in (2.4) the constraint $\mathbf{c}_{\beta,\ell_{N}}$ by $\mathbf{b}_{n,\ell_{N}}$, with
@@
In [17, Proposition 2.7] it is proved that the generator $\mathcal{B}^{n,\ell_{N}}$ is associated with a gradient model, in the sense of [21, II, Subsection 2.4]. Thus, from the linear combination in (2.8), our main process inherits this property. This is the content of the next lemma, whose proof is direct given the expression in [17, Proposition 2.7].
@@
Let $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{J}_{\beta}$}} be the algebraic current, defined through $\mathcal{L}^{\beta,\ell_{N}}[\pi_{0}]=\nabla\tau^{-1}[-\mathbf{J}_{\beta,\ell_{N}}]$, where $\mathbf{J}_{\beta,\ell_{N}}$ is given, explicitly, by
$\mathbf{J}_{\beta,\ell_{N}}=\frac{1}{\ell_{N}+1}\sum_{j=1}^{\ell_{N}}\mathbf{J}_{\beta}^{j}(\eta)\quad\text{with}\quad\mathbf{J}_{\beta}^{j}(\eta)=\beta(\langle\eta\rangle_{j})(\mathbf{e}_{0,1}(\eta)-\mathbf{e}_{1,0}(\eta)).$ (2.10)
@@
$\mathbf{J}_{\beta,\ell_{N}}=-\nabla\mathbf{H}_{\beta,\ell_{N}},$ (2.11)
@@
$\mathbf{g}_{\beta,\ell_{N}}(\eta)$ $=\frac{1}{\ell_{N}+1}\sum_{j=1}^{\ell_{N}}\sum_{i=1}^{j}\tau^{-i}\mathbf{J}_{\beta,\ell_{N}}^{j}(\eta),$ (2.13)
Remark 2.5.
It will be convenient to note an equivalent expression for $\mathbf{h}_{\beta,\ell_{N}}$. Let ${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{P}_{n}}}(A)$ be the collection of subsets of some set $A\subset\mathbb{T}_{N}$ with exactly $n$ elements. One can then express (see the last two displays in the proof of [17, Proposition 2.7] for the precise computations)
@@ 2.2. Topological setting
The uniqueness of solutions of the weak formulation in last definition is consequence of the regularity in (1). This is very straightforwardly shown via Oleinik’s method (see, for instance, [12, Lemma B.3]).
@@ 3.1. General scheme
The link between our process and the weak formulation (2.17) is given through Dynkin’s martingale (see [15, Appendix 1, Lemma 5.1]),
@@
One can show that the sequence of probability measures $(\mathbb{Q}_{N})_{N\in\mathbb{N}}$ is tight with respect to the Skorokhod topology of $\mathcal{D}\left([0,T],\mathcal{M}_{+}\right)$ by following [18, Subsection 3.2], which invokes Aldous’ conditions. As evidenced in [12, Proposition 3.1] and [18, Subsection 3.2], tightness for exclusion processes in the torus satisfying the gradient condition is consequence of the following: (i) $|\mathbf{r}_{\beta,\ell_{N}}|_{\infty}=o(N)$ ; (ii) $|\mathbf{h}_{\beta,\ell_{N}}|_{\infty}$ being uniformly bounded from above by a constant independent of $N$; (iii) and the estimate in Lemma 4.1 vanishing as $N\to\infty$. The former two are straightforwardly identified, therefore we refer the reader to the aforementioned references for the specific details to include them in the proof.
@@
Provided the aforementioned convergence by subsequences and absolutely continuity of the limit measure, one then argues that the density $\rho$ satisfies the notion of weak solution in Definition 2.6. This encompasses two results to be held $\mathbb{Q}$ almost-surely: (i) satisfaction of the weak formulation $\mathfrak{F}_{t}(\rho^{\text{ini}},\rho,G)=0$; and (ii) regularity, as in (1) in Definition 2.6. The latter is typically shown via the so-called ”energy estimate”. This is standard in the literature, and shown to hold for models with properties satisfied by the model in the present text. Concretely, proved the ”replacement lemmas” (here stated in Corollary A.2 and A.3), and the vanishing, as $N\to\infty$, of the remainder term in the gradient property, (2.13), as in Lemma 4.1, the regularity property (1) is consequence of $\mathbf{c}_{\beta,\ell_{N}}$ being uniformly bounded from above by a constant independent of $N$. For the precise computations, we refer the reader to [12, Section 5], where this was shown for the ”Interpolating process” in the slow-diffusion regime; or [18, Section 5 – Regime I]. The latter is closer, in terms of notation and exposition, to the present text. The next step is to characterize the limit $\mathbb{Q}$.
@@ 3.2. Characterization of the limiting points
where we recall that $\mathfrak{F}_{t}(\rho^{\rm ini},\rho,G)$ is given in (2.17). The next steps up to (3.14) are presented in detail in [18, Proof of Proposition 3.2], from the start of the proof up to equation (3.12) in the aforementioned work. Here, we only overview the approach and present the main novelties.
@@
  • After an application of Doob’s inequality, one can see that the first term vanishes in the limit $N\to+\infty$ by repeating the computations to show the first equality in [18, Equation (3.2)], where $|\mathbf{r}_{\beta,\ell_{N}}|_{\infty}=o(N)$ is required;
  • For the second, one applies [18, Proposition 4.3] and then Lemma 4.1;
@@
$\mathbf{h}_{\beta,\ell_{N}}(\eta)-\mathbf{h}_{\beta,L}(\eta)$ $=\sum_{m=0}^{M}\beta(\tfrac{m}{M})\big\{\frac{1}{\ell_{N}+1}\sum_{n=0}^{P_{\ell_{N}}(\eta)-1}\text{B}_{m,M}(\tfrac{n}{\ell_{N}})-\frac{1}{L+1}\sum_{n=0}^{P_{L}(\eta)-1}\text{B}_{m,M}(\tfrac{n}{L})\big\}$ (3.18)
@@
Applying [18, Proposition 4.3] and the triangle inequality, the limit above is a direct consequence of forthcoming Lemma 4.2.
@@
Next, we apply [18, Proposition 4.3] to each of the remaining probabilities and analyse them separately with the Replacement Lemmas in Section 4 (Corollaries A.2 and A.3), after an application of the triangle inequality in order to pass the summations over $m$ to outside of the expectation. Precisely, the terms associated with (3.28) and (3.27) are analysed with the upcoming Lemma A.2, with $L$ in the statement of the aforementioned lemma either equal to $1$ or an arbitrary integer $L>1$, respectively, and $w$ and $y$ chosen accordingly. To treat the term associated with (3.26), it is enough to apply the triangle inequality and then invoke Lemma A.3. In order to conclude the proof, we perform the sequence of limits $\limsup_{M\to+\infty}\limsup_{L\to+\infty}\limsup_{\epsilon\to 0}\limsup_{N\to+\infty}$.